OpenStudy (anonymous):

I do award medals!!! Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4. f(x) = −16x^2 f(x) = 16x^2 f(x) = −1/16 x^2 f(x) = 1/16 x^2

OpenStudy (anonymous):

I am thinking C

OpenStudy (anonymous):

f(x) = −1/16 x^2

OpenStudy (anonymous):

 \[\sqrt{\left( x-0 \right)^{2}+\left( y+4 \right)^{2}}=\frac{ \left| y-4 \right| }{ \sqrt{0^{2}+1^{2}} }\] squaring both sides \[x ^{2}+y ^{2}+8y+16=y ^{2}-8y+16\] \[x ^{2}=-16 y \] which is the required equation.